# Algorithms Trained on Normal Chest X-rays Can Predict Health Insurance Types

### v3

@@ 4 Discussion
Despite our results in AI algorithms predicting insurance information from chest X-ray images, there are several limitations in our work. First, we only chose one of the insurance types per person in their lifetime, which ignored the possible socioeconomic transition of a patient. Second, we did not explicitly show that health insurance information biases disease prediction in deep vision training, but since humans could easily learn the association between socioeconomic status and disease distribution, we should assume it being a simple association for a model to learn from. Third, we admit that the hyperparameters may not be best-tuned in the vision models due to computational limits, especially for the suspicious drop in DenseNet performance in isolated White people in MIMIC-CXR-JPG.
@@ 5 Conclusion
We demonstrated that the state-of-the-art deep vision models learn health insurance type information from chest X-ray images, and that it is not mediated by common demographic factors. Despite the convenience of AI technology, we should utilize these deep learning models more cautiously when applying them to real-world medical tasks.
@@ Appendix A Dataset description
MIMIC-IV v3.0 mimic1; mimic2 is a large medical dataset containing over 265,000 patients’ data collected at Beth Israel Deaconess Medical Center in Boston, MA, in the intensive care unit or emergency department between 2008-2022, while MIMIC-CXR-JPG mimic-cxr-jpg1; mimic-cxr-jpg2 is an extended image dataset for MIMIC-IV v3.0, including 377,110 chest X-ray images in total.
On the other hand, CheXpert chexpert includes 224,316 chest X-rays from 65,240 patients from both in-patient and out-patient centers of Stanford Hospital between October 2002 and July 2017. The recent CheXpert Plus paper chambon2024chexpert provides additional demographic information of each patient, including their health insurance type, race, sex, and age.
Appendix B Model architecture
DenseNet121 densenet comprises multiple convolutional neural network layers with denser residual connections than the traditional ResNet, which demonstrates equal importance of density with depth and width of a model. SwinTransformer swin is a variant of vision transformer model, where self-attention is performed on shifted image windows. This idea largely decreases the memory complexity and reduces the downsides of separated image areas. In our study, the version of SwinV2-B* was utilized specifically. Lastly, MedMamba medmamba combines both a convolutional branch and a state-space sequence model branch, Mamba, to utilize the local inductive bias and global feature extraction from each branch. As for the classification, a single linear layer was added after the penultimate layer of DenseNet121, SwinV2-B* and MedMamba.
All of the experiments utilized Lion lion as their optimizer, and the learning rate was 5e-6 for DenseNet, 1e-5 for SwinTransformer and MedMamba. The learning rate of DenseNet was adjusted to 1e-5 in experiment 3.3 for better performance. The batch size was 32 across our experiments, and we did not implement dropout or weight decay for regularization for simplicity. In MIMIC-CXR-JPG, all images were resized to 448x448 otherwise mentioned, while in CheXpert, all images were resized to 320x320 since the original image size in CheXpert is smaller than in MIMIC-CXR-JPG. All images were preprocessed with RandomHorizontalFlip, RandomRotation and Normalization in PyTorch for augmentation. A cross-entropy loss was utilized for health insurance type classification tasks. We trained our model on the training dataset, and selected the weights by the best performing average area under receiver operating characteristic curve (AUC) on the validation dataset. The final performance was tested on the testing dataset with the chosen weights. All experiments were trained on a single NVIDIA GeForce RTX 3090 or RTX 6000 Ada GPU.

### v4

@@ 4 Discussion
Interpreting these results requires considering what kinds of features AI models can extract from medical images. A useful conceptual framing distinguishes three types of features that medical-AI models can exploit: (1) biological features, directly reflecting anatomy or pathology; (2) social features, deriving from social determinants of health such as socioeconomic status or insurance types; and (3) hybrid features that intertwine both domains; for instance, age or smoking status, which combine biological and social dimensions. The ethical and technical goal is to minimize reliance on purely social features that reflect structural disadvantage rather than biology. Simply removing explicit demographic variables is insufficient as models often reconstruct them indirectly from other data. Fairness in medical AI therefore depends not only on data curation but on recognizing and attenuating hidden social signals embedded in the pixels themselves.
Our study design makes this challenge unusually visible. We trained and evaluated models solely on radiographs labeled “no finding,” images that lacked any CheXpert/MIMIC-defined abnormalities (cardiomegaly, enlarged cardiomediastinum, pulmonary edema, consolidation, pneumonia, pleural effusion, pneumothorax, atelectasis, fracture, lung lesion or opacity, and pleural or device-related findings), ensuring that all images were radiographically normal, frontal in view, and from patients under 65 years old. This design precludes shortcut learning from obvious pathology: the model could not rely on disease-related cues that might correlate with insurance status, yet its performance remained well above chance. That persistence implies the model captured either minute physiologic correlates of social conditions or contextual artifacts introduced during image acquisition. Some of these contextual artifacts may arise from system-level differences (such as hospital site, imaging hardware, or acquisition protocols) that correlate with insurance coverage, meaning the model could be detecting institutional rather than biological patterns. However, our experiments showed the strongest results using the single-site MIMIC dataset, indicating that other factors contribute to this phenomenon. The remaining signal likely originates in patient-level physiology itself, shaped over time by social environment, nutrition, stress, and comorbidity.
To examine where such subtle information might reside, we turned to spatial localization maps. Predictive regions were concentrated in the upper and mid-thorax, with minimal contribution from the lung bases. This pattern is unlikely to arise from artifacts such as tubes, leads, or devices, and may instead reflect subtle physiologic correlates. The upper and mid-thorax encompass the heart, great vessels, ribs, spine, and soft tissues; structures that may encode subtle, population-level differences linked to socioeconomic context. Variations in bone density, soft-tissue distribution, or vascular morphology shaped by chronic stress, nutrition, and healthcare access could all leave faint but detectable traces. The model’s focus on these regions aligns with prior work suggesting that AI systems can infer age, sex, and even mortality risk from texture and posture cues invisible to the human eye Ieki et al. (2022); Raghu et al. (2021); Li et al. (2022). Nevertheless, imaging-acquisition factors (e.g., variations in positioning, exposure, or calibration that correlate with insurance type) remain plausible contributors. Disentangling these biological and contextual pathways will require controlled studies linking quantitative anatomic features and imaging parameters to socioeconomic indicators.
If these cues are truly physiologic or contextual rather than demographic, they should persist across subgroups. Indeed, in experiments restricted to White participants, performance remained nearly unchanged. In another experiment predicting health insurance solely from combined demographic features also revealed low performance. These results suggest that demographic features is not acting as a direct proxy for insurance type in Chest X-rays. The model detects something more subtle; latent, physiologically mediated imprints of social conditions that escape coarse demographic categorization. This finding strengthens our argument that “invisible social fingerprints” can persist even when explicit social variables are absent, challenging the assumption that models trained on ostensibly neutral data are free from social bias.
@@
Despite our results in AI algorithms predicting insurance information from chest X-ray images, there are several limitations in our work. First, both the MIMIC-IV cohort we used and the current CheXpert-Plus dataset (as of February 2024) simplify a patient’s history by reporting only one insurance type. In MIMIC-IV, we specifically selected the insurance type associated with the lowest socioeconomic status. Both approaches fail to account for possible socioeconomic transitions during a patient’s lifetime. However, an analysis of the original MIMIC-IV dataset revealed that only around 8% of patients ever changed their insurance status, suggesting a limited impact by this transition factor. Second, we did not explicitly show that health insurance information biases disease prediction in deep vision training, but since humans could easily learn the association between socioeconomic status and disease distribution, we should assume it being a simple association for a model to learn from. Based on past literature discovering social biases in medical image predictions Seyyed-Kalantari et al. (2021), this is a reasonable assumption. Third, we admit that the hyperparameters may not be best-tuned in the vision models due to computational limits, especially for the suspicious drop in SwinV2-B* and MedMamba performance in isolated White people in CheXpert.
@@ 5 Conclusion
We demonstrated that the state-of-the-art deep vision models learn health insurance type information from chest X-ray images, and that it is not essentially mediated by common demographic factors. Despite the convenience of AI technology, we should utilize these deep learning models more cautiously when applying them to real-world medical tasks.
@@ Appendix A Dataset description
MIMIC-IV v3.0 Johnson et al. (2024, 2023) is a large medical dataset containing over 265,000 patients’ data collected at Beth Israel Deaconess Medical Center in Boston, MA, in the intensive care unit or emergency department between 2008-2022, while MIMIC-CXR-JPG Johnson et al. (2019a, b) is an extended image dataset for MIMIC-IV v3.0, including 377,110 chest X-ray images in total.
On the other hand, CheXpert Irvin et al. (2019) includes 224,316 chest X-rays from 65,240 patients from both in-patient and out-patient centers of Stanford Hospital between October 2002 and July 2017. The images were downsized to 390 x 320 in the downsized version. The recent CheXpert Plus paper Chambon et al. (2024) provides additional demographic information of each patient, including their health insurance type, race, sex, and age.